You are here: CLASSE Wiki>ACC/Bunch Web>FLMLog>FLM03WLog_2007 (07 Sep 2007, MarkPalmer)Edit Attach
Tags

Fast Luminosity Monitor Electronic Logbook 2007

23 Jul 2024 - 17:39

May 2007

May 9, 2007 T. Wilksen

Checked newly installed Motorola Coldfire module (Ethernet-based read out) for correct configuration settings of the boot environment and flashed ROM with CBI_NET version 0.96b.

Charlie Strohman replaced the p/s for the FLM03W module after the Coldfire did not boot up correctly before flashing it.

July 2007

July 16, 2007

Refurbished FLM detector installed in West Flare (M. Cole, G. Codner)

July 17, 2007

During machine studies shift, diagnosed lack of signals from detector. Traced to the fact that the readout module was powered due to an incompletely seated connector. Also, debugged a cabling mistake (vertical channels are read out via the Hi-Gain set of readouts in the module but were connected to the Low-Gain set of channels). Analog output found to be operational although gain calibrations are clearly needed. (M. Palmer, M. Cole, C. Strohman)

July 22, 2007 (M. Palmer, M. Cole)

Start detailed checkout and calibration using digital system.
  1. Review of FLM digital readout codes/procedures from EPT
    • Working directory (working on CESR28 for MPM access): u:[cesr.palmer.ts101.flm]
    • Run flm03w_reset.com
      1. Find that this has the WRONG polarity (first turns PS on and then off)
      2. Fix polarity in com file and put in additional status message
    • Run dl_flm.com
      1. Download of dsp_flm_v1.ldr completed successfully
      2. DSP was re-enabled automatically
    • Set up custom flm_v2.ini
      1. Copy flm_v2.ini to flm_v2_std.ini
      2. Copy flm_v2_mca.ini to flm_v2.ini and modify for only 20 data loops
    • Run flm_v2
      1. Initially get a crash (possibly because not hooked to a display)
      2. Set up display and repeat - program starts up successfully
  2. Set up for MCA (multi-channel analyzer data acquisition)
    • Short test run due to JTH tuning
      1. Should give rough calibration
      2. Specify 150 loops with bunches 3, 4, 5 selected
        • Acquisition Start: 20:23:57
        • Acquisition Complete: ~20:27:25
        • Data Upload (Ser. XBus) Complete: 20:29:52
        • Will need approx. 500 loops for proper calibration
        • Rename data file (default is mca_bunches.dat) to: mca_bunches_2007_0722_2023.dat [OOPS, this was the wrong file - contains bunch patterns, not MCA data].
    • Calibration Run
      1. Specify 500 loops (9 trains, bunches 3-5 only)
      2. Acquisition Start: 21:05:47
      3. Acquisition Complete: 21:16:20
      4. Data Upload Complete: 21:18:56
      5. Copy data files to:
        • flm_paw_MCA_2007_0722_2105.dat
        • flm_paw_nbin_2007_0722_2105.dat
      6. Transfer to Unix for analysis...

July 24, 2007 (M. Palmer, M. Cole, R. Mitchell)

  1. Set up for MCA data acquisition with horizontal angle (magnetic) through IP adjusted so that the horizontal scintillators near zero angle are illuminated.
    • Vilden puts in a regular fill
    • Adjust HBLIMING 2 (horizontal angle at L0) from 69 to 511 so that horizontal scintillators near zero crossing angle are being illuminated for calibration
  2. Start MCA data acquisition
    • Specify 750 loops (9 trains, bunches, 3-5 only)
    • Acquisition Start: 07:35:46
    • Acquisition Complete: 07:51:29
    • Data Upload Complete: 07:54:03
    • Copy data files to:
      • flm_paw_MCA_2007_0724_0735.dat
      • flm_paw_nbin_2007_0724_0735.dat
    • Transfer to Unix for analysis...

July 27, 2007 (M. Palmer, M. Cole)

  1. Clean up code needed for calibration sequence
    • Fix routine to calibrate pedestals so that it produces corrected pedestal values needed for constants file
      • [cesr.flm.do_gains]load_peds.f77
    • Correct pedestals taken on 7/25 and load. New pedestal values entered in flm_ped_gain.dat file in [cesr.flm.config]
    • Still see signs of saturation in vertical channels, may be due to operating at too high a voltage and entering the non-linear regime of the PMTs
    • Lower vertical voltages by 10 volts and see some variation across channels. Opt to run a new MCA calibration at this point.
    • Problems running the MCA routine so run flm03w_reset.com routine and lose all communications with DSP.
      • Turns out that there was a remaining typo in the com file and power to the readout module was still not getting enabled. Now fixed.
  2. MCA run (750 loops, 9 trains, bunches 3-5)
    • Acquisition Start: 16:44:33
    • Acquisition Complete: 17:00:17
    • Data Upload Complete: 17:02:34
    • Analyzed and installed new voltage settings ([cesr.flm.com]adjust_hv_2007_0728_1.com)
  3. 2nd MCA run with new voltages installed (500 loops, 9 trains, bunches 3-5)
    • Acquisition Start: 17:29:47
    • Acquisition Complete: 17:40:20
    • Data Upload Complete: 17:42:36
    • Transferred and analyzed data
    • Installed new voltage settings in [cesr.flm.com]adjust_hv_2007_0728_2.com
  4. 3rd MCA run with latest voltages (500 loops, 9 trains, bunches 3-5)
    • Acquisition Start: 18:00:25
    • Acquisition Complete: 18:10:58
    • Data Upload Complete: 18:13:15
    • Transferred and analyzed data

Sept. 7, 2007 (M. Palmer, M. Cole)

  1. Disable digital FLM acquisition
    • By running [cesr.palmer.ts101.flm]dca.com
  2. Start MCA run
    • copy flm_v2_mca.ini flm_v2.ini
    • Run flm_v2
    • Turn off unused bunches
    • Restarted after getting disconnected - going to look at program output and evaluate
    • Copy flm_paw_MCA.dat;1 flm_paw_mca_2007_0907.dat
    • Copy flm_paw_nbin.dat;1 flm_paw_nbin_2007_0907.dat
    • Created ADJUST_HV_2007_0907.COM and lowered HV by 25V for all vertical channels (b/c of signs of saturation in analog signals)
    • Applied new voltage settings
    • Compared analog and digital output for vertical channels and find good channel to channel agreement to within an overall scale factor
    • Edited [cesr.flm.config]flm_ped_gain.dat to adjust the vertical FLM analog channel gains
    • Downloaded new gain settings by running put_ped_gain.exe in same directory
    • Now see good agreement between analog and gdigital channel outputs
    • Will need to adjust the analog L/I calibration constant (but have to find it first!)
    • Disable digital readout again and take another set of calibration data
Topic revision: r8 - 07 Sep 2007, MarkPalmer
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding CLASSE Wiki? Send feedback